# Multi-modal Auto-regressive Modeling via Visual Words

### v1

@@ 3.1.1 Settings
In the specific implementation, we use CLIP-ViT-L-336px Radford et al. (2021) to initialise the visual encoder. for the multi-modal decoder and MM head, we use two different LLM initialisation schemes: the Vicuna-7B Vicuna (2023) and the Mistral-7B Jiang et al. (2023). VM head is a randomly initialised unbiased linear layer and adapter is a randomly initialised two-layer MLP. Please refer to appendix A for more detailed settings.
3.1.2 Stage I
At this stage, we aim to perform the preliminary alignment of the visual information into the semantic space of LLM. To achieve this, we employ adapter to project the uni-modal features output from the visual encoder into the semantic space of the multi-modal decoder, which serves as a contextual reference for the generation task. During this period, VW-LLM is trained for generating image caption. We train only adapter at this stage.
3.1.3 Stage II
At this stage, our goal is to adapt the LMM using multi-modal instruction data to obtain an LMM with multi-modal comprehension.At this stage, we train using more complex multi-modal instructions and a variety of tasks including complex inference, detailed description, multi-round dialogue, image description, and VQA. We train multi-modal decoder, MM head and adapter in this phase.
3.1.4 Stage III
At this stage, our goal is to train the VM head using an existing LMM and massive image data.We directly use the LMM to encode various images and use the final output of the MM head as the label.The VM head is then trained to fit the output of the MM head. We use KL divergence to compute the loss and only train the VM head at this stage.
3.1.5 Stage IV
In this stage, our goal is to use the VM head to construct the visual supervision information so as to train the LMM with multi-modal auto-regressive modelling. In this stage, we use the same multi-modal instruction dataset as stage II, but both language and visual information are involved in supervision in this stage of training. In this stage, we only train the multi-modal decoder and MM head.
3.2 Main Results
To evaluate the effectiveness of our proposed method, we conducted experiments on five widely used multi-modal benchmarks and four LMM benchmark toolkits. The experimental results are shown in Table 1. The experimental results demonstrate that our proposed VW-LMM, by constructing visual modeling tasks consistent with the pre-training tasks format of LLM, guides model to learn both language and visual modalities of the multi-modal sequences in an auto-regressive manner. This further bridges the gap between the two modal features and significantly enhances the model’s vision-language understanding capability.
Compared to models of the same scale, VW-LMM demonstrates superior or competitive performance on all evaluation metrics for vision-language understanding. Compared to larger-scale models, VW-LMM, with only the scale of 7B, outperforms some models with scale of 13B and achieves performance similar to models with a parameter scale of 33B. This once again verifies that our method, by introducing visual information supervision, further boosts the multi-modal understanding potential of LMM.
Compared to methods that solely apply language modeling as the optimization task, VW-LMM achieves significantly superior results, highlighting the importance of introducing visual insformation as supervision for LMMs. Existing methods that also employ multi-modal modeling tasks and perceive the supervision of visual information can be broadly classified into two categories. One involves constructing an additional image decoder to endow the model with image generation capabilities, thereby utilizing image information in the form of image denoising tasks. The other models the image information through regression tasks, aiming to directly fit the value of visual features using MSE loss. In comparison to the above methods, VW-LMM achieves better results by constructing visual words as supervision information, enabling a relatively unified and concise implementation of multi-modal modeling in the form of a classification task. This approach proves to be more effective and efficient.
It should be noted that both LaVIT and VW-LLM employ classification task for multi-modal modeling. However, LaVIT introduces an additional visual tokenizer and 16,384 trainable discrete embeddings. In contrast, Visual words utilize the existing semantic space of the pre-trained LLM to represent visual information, avoiding the introduction of massive training parameters. This approach also bridges the semantic gap between the two modalities and thus achieves better results.
3.3 Ablation
Methods LLM Res. Visual Question Answering Benchmark Toolkit
[…]
VW-LMM Vicuna-7B 336 78.9 62.7 48.3 68.1 57.6 85.9 65.9 59.8 31.3
VW-LMM Mistral-7B 336 80.8 65.4 58.5 75.9 63.1 87.0 80.6 79.0 44.0
VW-LLM with pseudo image features Vicuna-7B 336 77.2 61.6 48.1 65.4 54.5 85.6 65.2 53.5 30.1
Table 2: Results of ablation study and discussion.
In order to investigate whether the introduction of visual words as visual supervision information directly improves the vision-language comprehension performance of LMMs, we conducted an ablation study, and the experimental results refer to  2. From table  2, it can be observed that the introduction of visual words leads to evident performance enhancement in both different LLM settings, which confirms that VW-LMM, by constructing visual words as visual supervision information, guides the model to learn the rich semantics in multi-modal sequences through both language modelling and visual modelling, therefore boosting model’s vision-language understanding capability. In addition, we note that VW-LMM-Vicuna-7B shows a slight performance degradation on both VisWiz and TextVQA benchmarks. This case is due to the fact that VisWiz involves the requirement of the corresponding format of the model and TextVQA provides a large number of image OCR texts as reference. In both tasks, the language semantic understanding and instruction following capabilities of the model are more demanded. And the introduction of the visual modelling task slightly diminishes the text-side capabilities of Vicuna-7B on both datasets. When using Mistral-7B as the LLM, its stronger baseline performance and pre-trained language semantic comprehension bridged this gap, which in turn led to a consistent performance improvement.
4 Discussion
4.1 Replace Visual Feature with Pseudo Image Features
To further explore the manifestation of visual features in the semantic space of the LMM, we utilize the visual words and text embeddings to construct pseudo image features and replace visual features in the input sequence. The results are shown in table 2. From the experimental results, it is clear that the model that receives pseudo image features as visual input can still achieve vision-language understanding capability close to the original model. This result confirms that the embedding layer of the pre-trained LLM essentially achieves full utilisation of the semantic space, thus features from image modalities can be represented by linear combinations of text embeddings to some extent. Moreover, it also shows that visual words successfully implement the transfer of visual features to the language semantic space, again verifying the soundness of our proposed approach. However, the difference in results suggests that the structure of the linear projection might have limitations for the projection of visual information into the language semantic space, and we will further investigate how to better accomplish this operation in our subsequent work.
4.2 Visualization of Visual Words
Figure 4: Image regions with the highest probability tokens in the visual words. Best viewed zoomed-in.
In order to verify whether the visual words learnt by VW-LMM can realistically reflect the image information, we take VW-LMM-Vicuna-7B as an example to explore. For each patch in the image, we select the token with the highest probability in its corresponding visual words, and compare the region of interest in the image with its visualisation result, visualization is shown in figure  4.
From the visualisation results, it can be observed that: First of all, visual words can intuitively perceive the low-level features of images such as colour information and object boundary contours. The model in region (5) identifies the red region of the parrot’s face; the distribution of visual words in regions (5), (6) (8) clearly distinguishes the boundaries of the foreground and the background; and the visual words in regions (1), (3) accurately identifies the demarcation line between the building and the tree.
Secondly, visual words are able to recognise higher level semantics such as specific object categories. For example, roof, building, and tree are successfully recognised in regions (1) and (3); flag and window are recognised in regions (2) and (4); hair and feather of parrot are recognised in regions (5), (6), and (8); in region (4), visual words even mark the "walking" action of the crew.
In addition, we found that for the relationship between foreground and background in the internal semantic space of LMM, the model prefers to predict the background as a comma rather than a space or other placeholder as we expected. This phenomenon reflects the fact that visual words are simultaneously characterised by both visual and textual information, i.e. the background is seen in the image as a separator between different foreground patches.
The above phenomena indicate that visual words successfully achieve the transformation of visual features to language semantic space, which verifies the feasibility and effectiveness of our method. Therefore, using visual words as visual supervision signals can indeed guide the model to perform auto-regressive modelling of visual information, thus enhancing the model’s multi-modal comprehension and inference capabilities.
5 Related Work
5.1 Vision-Language Pre-training
Vision-Language Pre-training aims to build diverse multi-modal contextual training methods, thus endowing models with stronger multi-modal comprehension. Existing work has performed extensive and thorough research on Vision-Language Pre-training (VLP). CLIP Radford et al. (2021) proposes to apply contrastive learning with both visual encoder and text encoder to learn generic cross-modal representations, which lays the foundation of the learning paradigm for VLP. With the rapid development of LLM, researchers begin to explore LLM applied to vision-language tasks. Flamingo Alayrac et al. (2022) uses interleaved text-image data for training models with open generative capabilities; BLIP-2 Li et al. (2023a) and InstructBLIP Dai et al. (2023) use contrastive learning and ML tasks to construct efficient visual resampler for LLM at a low price. KOSMOS Huang et al. (2023) proposed to align the visual perception and language from scratch on web-scale multi-modal corpora. LLaVA series Liu et al. (2023b, a) construct complex Multi-modal instruction-following data and significantly improves model’s visual comprehension, and can be effectively extended to other vision-language tasks such as text understanding and region dialogue.
5.2 Visual Supervised Information
To further leverage visual information for supervised learning, Emu Sun et al. (2023b) and Emu2 Sun et al. (2023a) align visual and language modelling with the objective of predicting the next visual or language token in an auto-regressive manner, and further explore video as a new source of interleaved image-text data. This unified modelling provides a general framework for multi-modal understanding and generation. LaVIT Jin et al. (2023) performs tokenization of image features with trainable codebooks and trains external word embeddings in LLM for visual modelling.
However, the above methods either treat visual modelling as regression task, which does not match the pre-training format of LLM, or need to introduce extensive external training parameters and complex structures. In contrast to these methods, our approach uses LLM’s own word embedding space to construct visual words as supervised signals, using the classification task for visual modelling and do not need to introduce additional visual embedding.
6 Conclusion
In this paper, we successfully perform multi-modal auto-regressive modeling with a unified objective for the first time. Specifically, we propose the concept of visual words which transforms visual features in multi-modal sequences into probability distributions over the vocabulary of pre-trained LLMs, thus constructing supervision label for visual modeling. We also verified the representation of visual information by visual words and the feasibility of using visual words together with pre-trained embeddings to represent visual information in the language semantic space. The results show that visual words successfully achieve vision2language semantic transformation and effectively enhance the vision-language comprehension of model.
7 Limitations
Mapping continuous visual information into discrete language semantic space leads to a certain degree of information loss. This is also evidenced by the slight performance degradation brought about by the use of pseudo image feature. Furthermore, the effect of the diversity of training images on the learning effectiveness of the VM head was not explored. Therefore, it is necessary to further explore a more appropriate structure for constructing visual words and to verify the learning effect of VM head using richer image data.
Appendix A Detailed Settings
A.1 Hyperparameters.
We show the training hyper-parameters for the four stages of training in Table 3.
@@ Appendix A Detailed Settings
Table 3: Training hyper-parameters for the four stages of training.
A.2 Data
We show the dataset used in training of VM-LMM in Table 4.
Usage Source #Sample
Stage I LLaVA 1.5-558k 558k
Stage II LLaVA 1.5-mix-665k 665k
Stage III Images of LCS-558K 558k
Stage IV LLaVA 1.5-mix-665k 665k
Table 4: Dataset used in training of VM-LMM
For the stage I, we use the pre-trained data of LLaVA 1.5-558k Liu et al. (2023a). For the stage II and stage IV, we use the multi-modal instrction following data LLaVA-mix-665k Liu et al. (2023a). For the stage III, we use 1LCS-558K: a subset of 558K image-text pairs from LAION-CC-SBU with BLIP captions, as used in LLaVA-Lightning series.

### v2

@@ 3.1.1. Settings
In the specific implementation, we use CLIP-ViT-L-336px (Radford et al., 2021) to initialise the visual encoder. for the multi-modal decoder and MM head, we use two different LLM initialisation schemes: the Vicuna-7B (Vicuna, 2023) and the Mistral-7B (Jiang et al., 2023). VM head is a randomly initialised unbiased linear layer and adapter is a randomly initialised two-layer MLP. For detailed dataset information and training parameter settings, please refer to table 2 and table 3.
@@
Table 2. Training hyper-parameters for the four stages of training.
Usage Source
Stage I LLaVA 1.5-558k
Stage II LLaVA 1.5-mix-665k
Stage III Images of LCS-558K
Stage IV LLaVA 1.5-mix-665k
Table 3. Dataset used in training of VT-LMM
3.1.2. Stage I
At this stage, we aim to perform the preliminary alignment of the visual information into the semantic space of LLM. To achieve this, we employ adapter to project the uni-modal features output from the visual encoder into the semantic space of the multi-modal decoder, which serves as a contextual reference for the generation task. During this period, training task of VT-LMM is generating image caption for given image. We only train adapter with $Loss_{\mathrm{LM}}$ at this stage.
3.1.3. Stage II
At this stage, our goal is to adapt the LMM using multi-modal instruction data to obtain an LMM with multi-modal comprehension. At this stage, training task of VT-LMM involves a mixture of tasks, including complex inference, detailed description, multi-round dialogue, image caption, and visual question answering. We train multi-modal decoder, MM head and adapter with $Loss_{\mathrm{LM}}$ in this phase.
3.1.4. Stage III
In this stage, our objective is to train the VM head using existing LMM and massive image data, enabling it to map visual information to the language semantic space. The training task in this stage is to fit the output of LMM receiving pure image information using the VM head. It is important to note that in this stage, $Q$ serves as the label and $P^{\prime}$ serves as the logits. i.e., the training loss used in this stage is:
(9) $Loss_{\mathrm{VM}}^{\prime}=\frac{1}{|S_{\mathrm{{VM}}}|}\sum_{n\in S_{\mathrm{{VM}}}}\left(\sum_{i=0}^{C-1}Q_{n}(i)\log(\frac{Q_{n}(i)}{P^{\prime}_{n}(i)})\right).$
We only train VM head with $Loss_{\mathrm{VM}}^{\prime}$ in this phase.
3.1.5. Stage IV
In this stage, our goal is to use the VM head to construct the visual supervision information so as to train the LMM with multi-modal auto-regressive modelling. In this stage, we use the same dataset and training tasks as in Stage 2, but both language and visual information are involved in supervision. We only train the multi-modal decoder and MM head with $Loss_{\mathrm{MM}}$ at this stage.
Methods LLM Res. Adversarial Popular Random
Acc F1-Score Yes Acc F1-Score Yes Acc F1-Score Yes
LLaVA-v1.5(Liu et al., 2023b) Vicuna-13B 336 85.5 84.5 43.3 87.4 86.2 41.4 88.0 87.1 41.7
LLaVA-v1.5(Liu et al., 2023b) Vicuna-7B 336 85.1 84.2 44.0 87.2 86.1 41.9 88.1 87.3 41.9
mPLUG-Owl (Ye et al., 2023) LLaMA-7B 224 82.4 81.6 45.2 85.5 84.3 42.1 86.3 85.3 42.3
Multimodal-GPT (Gong et al., 2023) LLaMA-7B 224 50.0 66.7 100.0 50.0 66.7 100.0 50.0 66.7 100.0
MoE-LLaVA-2.7B×4-Top2 (Lin et al., 2024) Phi-2-2.7B 336 85.9 84.9 43.2 87.5 86.4 41.8 88.5 87.7 41.8
VT-LMM Vicuna-7B 336 85.1 84.1 44.0 87.4 86.3 41.7 88.1 87.3 41.9
VT-LMM Mistral-7B 336 87.3 86.1 41.9 88.3 87.1 40.9 88.7 87.8 41.4
Table 4. Results of object hallucination evaluation. Yes denotes the proportion of answering “Yes” to the given question. The best results and second best results are indicated by boldface and underline, respectively.
Methods LLM Res. Visual Question Answering Benchmark Toolkit
[…]
VT-LMM Vicuna-7B 336 78.9 62.7 48.3 68.1 57.6 85.9 65.9 59.8 31.3
VT-LMM Mistral-7B 336 80.8 65.4 58.5 75.9 63.1 87.0 80.6 79.0 44.0
VT-LMM with pseudo image features Vicuna-7B 336 77.2 61.6 48.1 65.4 54.5 85.6 65.2 53.5 30.1
Table 5. Results of ablation study and discussion.
3.2. Main Results
To evaluate the effectiveness of our proposed method, we conducted experiments on five widely used multi-modal benchmarks and four LMM benchmark toolkits. The experimental results are shown in Table 1. The experimental results demonstrate that our proposed VT-LMM, by constructing visual modeling tasks consistent with the pre-training tasks format of LLM, guides model to learn both language and visual modalities of the multi-modal sequences in an auto-regressive manner. This further bridges the gap between the two modal features and significantly enhances the model’s vision-language understanding capability.
Compared to models of the same scale, VT-LMM demonstrates superior or competitive performance on all evaluation metrics for vision-language understanding. Compared to larger-scale models, VT-LMM, with only the scale of 7B, outperforms some models with scale of 13B and achieves performance similar to models with a parameter scale of 33B. This once again verifies that our method, by introducing visual information supervision, further boosts the multi-modal understanding potential of LMM.
Compared to methods that solely apply language modeling as the optimization task, VT-LMM achieves significantly superior results, highlighting the importance of introducing visual insformation as supervision for LMMs. Existing methods that also employ multi-modal modeling tasks and perceive the supervision of visual information can be broadly classified into two categories. One involves constructing an additional image decoder to endow the model with image generation capabilities, thereby utilizing image information in the form of image denoising tasks. The other models the image information through regression tasks, aiming to directly fit the value of visual features using MSE loss. In comparison to the above methods, VT-LMM achieves better results by constructing visual tokens as supervision information, enabling a relatively unified and concise implementation of multi-modal modeling in the form of a classification task. This approach proves to be more effective and efficient.
It should be noted that both LaVIT and VT-LMM employ classification task for multi-modal modeling. However, LaVIT introduces an additional visual tokenizer and 16,384 trainable discrete embeddings. In contrast, visual tokens utilize the existing semantic space of the pre-trained LLM to represent visual information, avoiding the introduction of massive training parameters. This approach also bridges the semantic gap between the two modalities and thus achieves better results.
3.3. Object Hallucination Evaluation
We adopt the Polling- based Object Probing Evaluation (POPE) to evaluate object hallucination in VT-LMM and report results in table 4. The two variants of VT-LMM have achieved excellent performance in models of the same scale. Notably, VT-LMM-Mistral-7B has achieved the best performance across three different data sampling methods: Adversarial, Popular, and Random. Additionally, the yes ratio of VT-LMM remains relatively balanced, indicating that our method can provide accurate identification results based on the given query. These results demonstrate that our VT-LMM, achieving unified multi-modal auto-regressive modeling across multi-modal sequences by introducing visual tokens, explicitly forces the model to learn the semantic distributions of both visual and language information, leading to improved visual-language consistency. i.e., The model tends to perceive and generate object information that aligns with the given image content.
Figure 4. Image regions with the highest probability tokens in the visual tokens. Best viewed zoomed-in.
3.4. Ablation
In order to investigate whether the introduction of visual tokens as visual supervision information directly improves the vision-language comprehension performance of LMMs, we conducted an ablation study, and the experimental results refer to table 5. From table  5, it can be observed that the introduction of visual tokens leads to evident performance enhancement in both different LLM settings, which confirms that VT-LMM, by constructing visual tokens as visual supervision information, guides the model to learn the rich semantics in multi-modal sequences through both language modelling and visual modelling, therefore boosting model’s vision-language understanding capability. In addition, we note that VT-LMM-Vicuna-7B shows a slight performance degradation on both VisWiz and TextVQA benchmarks. This case is due to the fact that VisWiz involves the requirement of the corresponding format of the model and TextVQA provides a large number of image OCR texts as reference. In both tasks, the language semantic understanding and instruction following capabilities of the model are more demanded. And the introduction of the visual modelling task slightly diminishes the text-side capabilities of Vicuna-7B on both datasets. When using Mistral-7B as the LLM, its stronger baseline performance and pre-trained language semantic comprehension bridged this gap, which in turn led to a consistent performance improvement.
4. Discussion
4.1. Replace Visual Feature with Pseudo Image Features
To further explore the manifestation of visual features in the semantic space of the LMM, we utilize the visual tokens and text embeddings to construct pseudo image features and replace visual features in the input sequence. The results are shown in table 5. From the experimental results, it is clear that the model that receives pseudo image features as visual input can still achieve vision-language understanding capability close to the original model. This result confirms that the embedding layer of the pre-trained LLM essentially achieves full utilisation of the semantic space, thus features from image modalities can be represented by linear combinations of text embeddings to some extent. Moreover, it also shows that visual tokens successfully implement the transfer of visual features to the language semantic space, again verifying the soundness of our proposed approach. However, the difference in results suggests that the structure of the linear projection might have limitations for the projection of visual information into the language semantic space, and we will further investigate how to better accomplish this operation in our subsequent work.
4.2. Visualization of Visual Tokens
In order to verify whether the visual tokens learnt by VT-LMM can realistically reflect the image information, we take VT-LMM-Vicuna-7B as an example to explore. For each patch in the image, we select the token with the highest probability in its corresponding visual tokens, and compare the region of interest in the image with its visualisation result, visualization is shown in figure  4.
From the visualisation results, it can be observed that: First of all, visual tokens can intuitively perceive the low-level features of images such as colour information and object boundary contours. The model in region (5) identifies the red region of the parrot’s face; the distribution of visual tokens in regions (5), (6) (8) clearly distinguishes the boundaries of the foreground and the background; and the visual tokens in regions (1), (3) accurately identifies the demarcation line between the building and the tree.
Secondly, visual tokens are able to recognise higher level semantics such as specific object categories. For example, roof, building, and tree are successfully recognised in regions (1) and (3); flag and window are recognised in regions (2) and (4); hair and feather of parrot are recognised in regions (5), (6), and (8); in region (4), visual tokens even mark the ”walking” action of the crew.
In addition, we found that for the relationship between foreground and background in the internal semantic space of LMM, the model prefers to predict the background as a comma rather than a space or other placeholder as we expected. This phenomenon reflects the fact that visual tokens are simultaneously characterised by both visual and textual information, i.e. the background is seen in the image as a separator between different foreground patches.
The above phenomena indicate that visual tokens successfully achieve the transformation of visual features to language semantic space, which verifies the feasibility and effectiveness of our method. Therefore, using visual tokens as visual supervision signals can indeed guide the model to perform auto-regressive modelling of visual information, thus enhancing the model’s multi-modal comprehension and inference capabilities.
5. Related Work
5.1. Vision-Language Pre-training
Vision-Language Pre-training aims to build diverse multi-modal contextual training methods, thus endowing models with stronger multi-modal comprehension. Existing work has performed extensive and thorough research on Vision-Language Pre-training (VLP). CLIP (Radford et al., 2021) proposes to apply contrastive learning with both visual encoder and text encoder to learn generic cross-modal representations, which lays the foundation of the learning paradigm for VLP. ALBEF (Li et al., 2021) further improves the learning effectiveness of VLP by applying Image-Text Contrastive loss (ITC) and Image-Text Matching task (ITM) to the classical Masked Language Modelling (MLM) task to further enhance the learning of cross-modal representations. BLIP (Li et al., 2022) combines ITM, ITC and language modelling task (LM) as a classical work on multi-modal generative training methods. With the rapid development of LLM, researchers begin to explore LLM applied to vision-language tasks. Flamingo (Alayrac et al., 2022) uses interleaved text-image data for training models with open generative capabilities; BLIP-2 (Li et al., 2023b) and InstructBLIP (Dai et al., 2023) use contrastive learning and ML tasks to construct efficient visual resampler for LLM at a low price. KOSMOS (Huang et al., 2023) proposed to align the visual perception and language from scratch on web-scale multi-modal corpora. Previous VLPs used image-text pairs datasets and simple VQA datasets with limited data diversity and task complexity to enhance the general understanding and generation of Large Multi-modal Model (LMM). With the progress of instruction-following in LLM, Multi-modal instruction-following task is proposed. LLaVA series (Liu et al., 2023c, b) construct complex Multi-modal instruction-following data and significantly improves model’s visual comprehension, and can be effectively extended to other vision-language tasks such as text understanding and region dialogue.
5.2. Visual Supervised Information
To further leverage visual information for supervised learning, Emu (Sun et al., 2023b) and Emu2 (Sun et al., 2023a) align visual and language modelling with the objective of predicting the next visual or language token in an auto-regressive manner, and further explore video as a new source of interleaved image-text data. This unified modelling provides a general framework for multi-modal understanding and generation. LaVIT (Jin et al., 2023) utilizes narest neighbor lookup to map image information to discrete trainable codebooks, and extends the original vocabulary of LLM to construct one-hot supervised labels for image information. Unified-IO 2 (Lu et al., 2023) improves the modeling capability of the backbone for multi-modal information by predicting corrupt modalities using an additional VQ-GAN decoder and optimizing the pixel-level reconstruction loss.
However, the above methods either treat visual modelling as regression task, which does not match the pre-training format of LLM, or need to introduce extensive external training parameters and complex structures. In contrast to these methods, our approach uses LLM’s own word embedding space to construct visual tokens as supervised signals, using the classification task for visual modelling and do not need to introduce additional visual embedding.
6. Conclusion
In this paper, we successfully perform multi-modal auto-regressive modeling with a unified objective for the first time. Specifically, we propose the concept of visual tokens which transforms visual features in multi-modal sequences into probability distributions over the vocabulary of pre-trained LLMs, thus constructing supervision label for visual modeling. We also verified the representation of visual information by visual tokens and the feasibility of using visual tokens together with pre-trained embeddings to represent visual information in the language semantic space. The results show that visual tokens successfully achieve vision2language semantic transformation and effectively enhance the vision-language comprehension of model.
7. Limitations
Mapping continuous visual information into discrete language semantic space leads to a certain degree of information loss. This is also evidenced by the slight performance degradation brought about by the use of pseudo image feature. Furthermore, the effect of the diversity of training images on the learning effectiveness of the VM head was not explored. Therefore, it is necessary to further explore a more appropriate structure for constructing visual tokens and to verify the learning effect of VM head using richer image data.
Acknowledgement
Corresponding author: Zuchao Li and Ping Wang. This work was supported by the National Natural Science Foundation of China (No. 62306216), the Natural Science Foundation of Hubei Province of China (No. 2023AFB816), the Fundamental Research Funds for the Central Universities (No. 2042023kf0133), the Special Fund of Hubei Luojia Laboratory (No. 220100014), National Natural Science Foundation of China [No. 72074171] [No. 72374161].